Find the weight of H, stored in 100-cm' cylindrical tank at 100 °C and 100 atm (10, 132.5 KPa) using (a) Ideal gas equation (b) van der Waals equation if a = 0.244 atm-L²/mole and b = 0.0266 L/mole. (Note: Use the method of successive approximation to solve for the mass of H₂)

1
Expert's answer
2022-02-28T11:02:36-0500

V= 100cm3= 0.1L

T= 373k

P= 100atm

R= 10÷132.5= 0.0755

n= ?

n= PV/RT

n= (0.1×100)÷ (0.0755×373)

= 0.355